Balance Sheet Summary for Vandelay Farms Historica Pro 2017 Current assets Cash, savings, and marketable securities Crops, feed, and supplies Market livestock Accounts receivable Other current assets All personal assets Total current assets 16,853 110,265 16,689 110,708 111,133 94,835 109,811 111,841 95,564 35,831 290,749 289,518 290,116 289,954 291,752 18,791 36,548 35,588 12,866 current 13,418 Noncurrent assets Machinery and equipment Farm real estate Other noncurrent assets All personal assets Total noncurrent assets Total assets 47,249 492,453 47,005 47,838 493,412 47,395 251,279 251,393 252,133 252,360 253,171 492,638 492,456 493,267 46,991 851,486 854,080 854,893 853,635 854,792 1,142,235 1,143,598 1,145,009 1,143,589 1,146,544 46,024 45,364 Liabilities (S) Current liabilities Operating, short-term notes Current maturities Accounts payable Other current liabilities Personal current liabilities Total current liabilities 94,123 25,897 8,456 84,963 800 94,950 26,648 95,409 96,278 8,724 85,969 9,420 86,665 85,857 214,239 216,959 217,286 218,879 219,624 Noncurrent liabilities Non-real estate farm debt Real estate farm debt Other noncurrent liabilities Personal noncurrent liabilities Total noncurrent liabilities Total liabilities 83,204 250,263 83,387 249,971 82,597 249,128 249,046 249,789 3,069 2,924 346,858 346,738 346,677 345,954 346,547 563,963 561,097 563,697 564,833 566,171 Contributed capital and retained earnings Valuation/personal asset equity Total owner equity Total liabilities and owner equity 346,584 234,462 581,046 1,142,235 1,143,598 1,145,009 1,143,589 1,146,544 346,987 346,070 234,151233,831 579,901 346,558 232,198233,20:1 578,756 347,172 581,138 Balance sheet notes Farm assets Farm liabilities Farm equity 1,080,971 1,082,785 1,085,152 1,084,612 1,088,218 561,041 527,177 560,634 560,516 523,455 523,870

Income Statement Summary for Vandelay Farms Historica Pro forma Item (S) Value of farm production Non-depreciation operating expenses Depreciation expense nterest expense Net income from operations Gain (loss) on sale of capital items Net farm income before taxes Nonfarm income Income before taxes Income and social security taxes Net income after taxes Income statement notes Interest and dividend income received on farm assets 2014 481,964 385,975 28,156 34,754 33,079 2015 484,387 383,177 24,346 35,989 40,875 2016 487,876 385,197 26,285 37,664 38,730 2017 488,064 381,853 27,237 35,585 43,389 2018 490,447 383,397 29,201 35,780 42,069 33,079 24,156 57,235 7,011 50,224 40,875 27,552 68,427 4,360 64,067 38,730 29,206 67,936 7,603 60,333 43,389 28,308 71,697 6,686 65,011 42,069 29,364 71,433 2,877 556 680 740 696 701

Statement of Cash Flow Summary for Vandelay Farms Pro Cash flows from activities 2015 Cash received from farm production and govt. payments Cash received from net nonfarm income Total 526,987 555,689 27,552 593,654 570,879 24,156 Cash paid for operating expenses Cash paid for interest on Cash paid for interest on term loans Cash paid for feed and market livestock Cash paid for ather items purchased for resale Cash paid for income and social security taxes Cash withdrawals for farnily living Total 372,923 8,123 374,836 8,129 20,063 376,927 372,552 8,323 8,380 19,221 82,521 19,976 83,456 6,107 56,987 547,159 4,009 58,030 7,506 60,739 2,965 60,847 543,842 59,911 Net cash flows provided by operating activities Cash flows from investing activities Cash received on sale of machinery,equipment and real estate Cash received from sale of breeding livestock Cash received from withdrawals of savings, marketable securities, personal Total Cash paid to purchase machinery, equipment, and real estate Cash paid for purchase of breeding livestock Cash paid for deposits to savings accounts, marketable securities, personal 6,158 26,879 115,389 2,500 20,159 47,038 47,038 22,798 138,187 Net cash flows from investing activities Cash flows from financing activities Proceeds from operating loans and short-term notes Proceeds from debt financing Cash received from capital contributions 6,145 115,648 789 15,789 Cash repayment of aperating and short-term loans Cash repayment of term debt and capital leases-scheduled Cash repayment of term debt and capital leases-unscheduled Cash payments of dividends and other capital distributions Total 4,178 13,789 26,123 39,753 753 Net cash flows from financing activities 22,766 92,817 Net increase (decrease) in cash flows Beginning of year cash balance End of year cash balance 11,031 18,154 29,185 9,561 12,780 15,350 12,780

Given the above Finacail Statemtns, Calculate all of the missing performance measures in sheet “PM” (below) for Vandelay Farms.“CF” is the statement of cash flow summary for Vandelay Farms. Based on the performance ratios, is Vandelay Farms in any kind of financial trouble? Why or why not?

Coverage and repayment capacity Interest coverage ratio Term debt and capital lease coverage ratio Replacement margin ratio Capital debt repayment capacity Capital debt repayment margin Replacement margin
Answer #1

Answer :

Interest Coverage ratio = EBIT/Interest expense

where EBIT = Earning before interest and tax

2014

Net income from operation = 33079  

Interest expense = 34754

EBIT = Net income from operation + Interest expense (33079 + 34754) = 67833

Interest Coverage ratio = EBIT/Interest expense = 67833/34754 = 1.95 approx.

Term debt and lease service coverage ratio = Net operating income / Total debt service

EBIT = Net income from operation + Interest expense (33079 + 34754) = 67833

Total debt = 561097

Term debt and lease service coverage ratio = 67833/561097 = 0.12 approx.

Capital Debt Repayment Capacity = Net Income + Depreciation Expense + Non-Farm/Business Income – Family Living Expenses & Income Taxes + Interest Expense on Term Loans=33079 + 28156 + 24156 – 7011 + 19563= 97943

2015

Net income from operation = 40875

Interest expense = 35989

EBIT = Net income from operation + Interest expense (40875 + 35989) = 76864

Interest Coverage ratio = EBIT/Interest expense = 76864/35989= 2.14 approx.

Term debt and lease service coverage ratio = Net operating income / Total debt service

EBIT = Net income from operation + Interest expense (40875 + 35989) = 76864

Total debt = 563697

Term debt and lease service coverage ratio = 76864/563697 = 0.13 approx.

Capital Debt Repayment Capacity = Net Income + Depreciation Expense + Non-Farm/Business Income – Family Living Expenses & Income Taxes + Interest Expense on Term Loans=40875+ 24346+ 27552– 4360+ 20063= 124402

Capital Debt Repayment Margin = Capital Debt Repayment Capacity – scheduled principal and interest on term loans and leases = 108476– (4178+ 20063) = 84235

Replacement margin ratio = Capital Debt Repayment Capacity / (scheduled loans and leases + cash use to purchase replace asset) = 108476/ (4178+ 20063+ 6158) = 3.568

2016

Net income from operation = 38730

Interest expense = 37664

EBIT = Net income from operation + Interest expense (38730 + 37664) = 76394

Interest Coverage ratio = EBIT/Interest expense = 76394/37664= 2.02 approx.

Term debt and lease service coverage ratio = Net operating income / Total debt service

EBIT = Net income from operation + Interest expense (38730 + 37664) = 76394

Total debt = 563963

Term debt and lease service coverage ratio = 76394/563963 = 0.135 approx.

Capital Debt Repayment Capacity = Net Income + Depreciation Expense + Non-Farm/Business Income – Family Living Expenses & Income Taxes + Interest Expense on Term Loans=38730+ 26285+ 29206– 7603+ 19976= 126570

Capital Debt Repayment Margin = Capital Debt Repayment Capacity – scheduled principal and interest on term loans and leases = 126570– (13789+ 19976) = 92805

Replacement margin ratio = Capital Debt Repayment Capacity / (scheduled loans and leases + cash use to purchase replace asset) = 126570/ (13789+ 19976+ 23489) = 1.995

2017

Net income from operation = 43389

Interest expense = 35585

EBIT = Net income from operation + Interest expense (43389 + 35585) = 78974

Interest Coverage ratio = EBIT/Interest expense = 78974/35585= 2.22 approx.

Term debt and lease service coverage ratio = Net operating income / Total debt service

EBIT = Net income from operation + Interest expense (43389 + 35585) = 78974

Total debt = 564833

Term debt and lease service coverage ratio = 78974/564833 = 0.139 approx.

Capital Debt Repayment Capacity = Net Income + Depreciation Expense + Non-Farm/Business Income – Family Living Expenses & Income Taxes + Interest Expense on Term Loans=43389+ 27237+ 28308– 6686 + 19221= 111469

Capital Debt Repayment Margin = Capital Debt Repayment Capacity – scheduled principal and interest on term loans and leases = 111469– (12432+ 19221) = 79816

Replacement margin ratio = Capital Debt Repayment Capacity / (scheduled loans and leases + cash use to purchase replace asset) = 111469/ (12432+ 19221+ 26879) = 1.904

2018

Net income from operation = 42069

Interest expense = 35780

EBIT = Net income from operation + Interest expense (42069 + 35780) = 77849

Interest Coverage ratio = EBIT/Interest expense = 77849/35780= 2.17 approx.

Term debt and lease service coverage ratio = Net operating income / Total debt service

EBIT = Net income from operation + Interest expense (42069 + 35780) = 77849

Total debt = 566171

Term debt and lease service coverage ratio = 77849/566171 = 0.137 approx.

Capital Debt Repayment Capacity = Net Income + Depreciation Expense + Non-Farm/Business Income – Family Living Expenses & Income Taxes + Interest Expense on Term Loans=42069 + 29201 + 29364 – 2877 + 19079= 116836

Replacement margin ratio = Capital Debt Repayment Capacity / (scheduled loans and leases + cash use to purchase replace asset) = 116836/ (0 + 115648) = 1.010

Conclusion :

Based on performance ratio, Vandelay firm is not in any kind of trouble, as over the years the debt service coverage ratio has increased. The financial position of the company is sound.